### v1

# Fine-Grained ECG-Text Contrastive Learning via Waveform Understanding Enhancement

Abstract
Electrocardiograms (ECGs) are essential for diagnosing cardiovascular diseases. While previous ECG-text contrastive learning methods have shown promising results, they often overlook the incompleteness of the reports. Given an ECG, the report is generated by first identifying key waveform features and then inferring the final diagnosis through these features. Despite their importance, these waveform features are often not recorded in the report as intermediate results. Aligning ECGs with such incomplete reports impedes the model’s ability to capture the ECG’s waveform features and limits its understanding of diagnostic reasoning based on those features. To address this, we propose FG-CLEP (Fine-Grained Contrastive Language ECG Pre-training), which aims to recover these waveform features from incomplete reports with the help of large language models (LLMs), under the challenges of hallucinations and the non-bijective relationship between waveform features and diagnoses. Additionally, considering the frequent false negatives due to the prevalence of common diagnoses in ECGs, we introduce a semantic similarity matrix to guide contrastive learning. Furthermore, we adopt a sigmoid-based loss function to accommodate the multi-label nature of ECG-related tasks. Experiments on six datasets demonstrate that FG-CLEP outperforms state-of-the-art methods in both zero-shot prediction and linear probing across these datasets.
Fine-Grained ECG-Text Contrastive Learning via Waveform Understanding Enhancement
Haitao Li1, Che Liu2, Zhengyao Ding1, Ziyi Liu3, Zhengxing Huang1 1Zhejiang University, 2Imperial College London, 3Transtek Medical Electronics Co., Ltd lihaitao@zju.edu.cn, zhengxinghuang@zju.edu.cn
1 Introductin
Electrocardiograms (ECGs) are important, non-invasive tools used in clinical settings to detect heart rhythm disorders Sahoo et al. (2020); Rath et al. (2021); Ayano et al. (2022). ECG self-supervised learning, which uses unlabeled data, has become a promising solution to the limitations of traditional methods that rely on annotated data and require expert knowledge. It can be broadly classified into two types: comparative self-supervision Chen et al. (2020, 2021); Wang et al. (2023); Eldele et al. (2021), which focuses on learning ECG features by distinguishing between augmented positive and negative samples, and generative self-supervision Zhang et al. (2022a); Hu et al. (2023); Na et al. (2024); Zhang et al. (2022b), which aims to reconstruct original ECG signals from masked versions. However, these methods often require annotated samples and struggle with classes not present in the fine-tuning stages.
Figure 1: (a) Doctors typically make diagnoses based on waveform features. (b) Using LLMs to recover waveform features from diagnosis faces two challenges: hallucinations and the non-bijective relationship
Significant efforts inspired by vision-text multimodal contrastive learning CLIP Radford et al. (2021) have been made to address such zero-shot challenges. Li et al. (2024); Liu et al. (2024a); Lalam et al. (2023) learn ECG representations by pulling embeddings of paired ECG signals and their corresponding reports while pushing them from unpaired reports, enabling zero-shot prediction. MERL Liu et al. (2024b) further enhances these representations with uni-modal alignment and employs the CKEPE pipeline during inference to generate more descriptive prompts using LLMs. However, enhancing textual prompts only at the inference stage can’t fully leverage the capabilities of LLMs. In contrast, ESI Yu et al. (2024) enhances ECG reports during training by integrating a retrieval-augmented generation (RAG) pipeline that leverages LLMs and external medical knowledge to produce detailed textual descriptions.
However, none of the aforementioned methods noted the missing waveform features in ECG reports. As shown in Figure 1(a), when given an ECG, doctors typically first recognize key waveform features and then make a final diagnosis based on these features. However, these intermediate waveform features are often not recorded in the report, despite being crucial for the model to understand the diagnostic process. Our objective is to recover these omitted features. Notably, simply augmenting the report using LLMs Yu et al. (2024), as shown in Figure 1(b), is unreliable due to: (1) the hallucination of medical LLMs Huang et al. (2023); Günay et al. (2024) and (2) the non-bijective relationship between ECG waveform features and diagnoses Jin (2018), where a single disease may present different waveform characteristics across individuals.
In this study, we propose a novel training strategy FG-CLEP to address the aforementioned challenge. It consists of the following three steps: (1) training the CLEP model using contrastive learning on original ECG-report pairs, (2) generating potential waveform features based on the original report using LLMs and validating them with CLEP, and (3) continuing to train the CLEP model with this augmented report containing waveform features to obtain the final FG-CLEP model. The key is to use the coarse CLEP model to calculate the similarity between the ECG signal and the potential waveform features generated by LLMs. High-probability features are treated as ground truth and incorporated into the reports, which are then used for FG-CLEP training. Thus, this approach not only addresses the non-bijective relationship problem but also provides a mechanism for error correction when LLMs exhibit hallucination issues.
Additionally, ECG data possess unique characteristics compared to images, making direct reuse of the CLIP architecture Radford et al. (2021) suboptimal. First, most ECGs are normal Thai et al. (2017); Yogarajan et al. (2021), leading to similar report semantics across patients. However, previous methods Radford et al. (2021); Liu et al. (2024b, a) treat only ECG-report pairs from the same patient as positive samples, causing frequent false negatives during pre-training. Thus, we introduced a semantic similarity matrix to measure the semantic similarity between different reports, enabling the identification and elimination of false negatives. Second, ECG tasks typically require multi-label classification, but prior approaches use a softmax-based Info-NCE loss Oord et al. (2018), which is unsuitable for such tasks. To address this, we replaced Info-NCE with a sigmoid-based loss to support multi-label scenarios.
Figure 2: Training Porcess of FG-CLEP. (1) training the CLEP model using contrastive learning on original ECG-report pairs, (2) generating potential waveform features based on the original report using LLMs and validating them with CLEP, and (3) continuing to train the CLEP model with this augmented report containing waveform features to obtain the final FG-CLEP model.
We validate our proposed FG-CLEP on six ECG multi-label classification datasets in both zero-shot and linear probing, demonstrating competitive performance compared to state-of-the-art methods in both settings. Overall, our contributions are threefold:
  • We propose a novel training pipeline to capture fine-grained waveform features. Since LLM-generated waveform features are unreliable due to hallucination and the non-bijective relationship, our pipeline validates these features by measuring their similarity to the ECG signal, generating a refined report for continued training.
  • We design a model tailored for ECG multi-label nature and frequent false negatives. We implement a sigmoid-based loss function and introduce a semantic similarity matrix to measure the similarity of reports across patients, guiding contrastive learning and effectively mitigating false negatives.
  • Experimental results indicate that the model, pre-trained on MIMIC-ECG using our FG-CLEP framework, outperforms state-of-the-art methods in both zero-shot prediction and linear probing across six datasets, including PTB-XL, CPSC2018, and CSN.
2 Related Work
ECG Self-Supervised Learning Self-supervised learning in ECG analysis has primarily been explored through two paradigms: Contrastive self-supervision Chen et al. (2020, 2021); Wang et al. (2023); Eldele et al. (2021), which typically involves augmenting the same ECG signal into two different views as positive samples, while different ECG signals serve as negative samples; Generative self-supervision Zhang et al. (2022b); Hu et al. (2023); Zhang et al. (2022a); Na et al. (2024), which first masks a portion of the ECG signal and then attempts to recover the masked part using the unmasked portion. Unlike previous ECG self-supervised methods that rely on annotations and struggle with unseen classes during fine-tuning, our FG-CLEP enables direct zero-shot prediction on downstream tasks.
ECG-Report Contrastive Learning Recently, inspired by the strong zero-shot ability of image-caption multimodal contrastive learning methods like CLIP Radford et al. (2021), significant efforts have been made in ECG-Report contrastive learningLi et al. (2024); Liu et al. (2024b, a); Yu et al. (2024); Lalam et al. (2023). Similar to CLIP Radford et al. (2021), Li et al. (2024); Liu et al. (2024a); Lalam et al. (2023) learns ECG representations by pulling ECGs with their paired reports while pushing them from unpaired reports. MERL Liu et al. (2024b) further introduces uni-modal alignment and employs the CKEPE pipeline at inference to generate more descriptive prompts via LLMs. However, enhancing textual prompts only during inference creates a distribution mismatch between training and testing text. In contrast, ESI Yu et al. (2024) enhances ECG reports during training using a retrieval-augmented generation (RAG) pipeline, integrating LLMs and external medical knowledge for more detailed descriptions.
Despite these advances, existing methods overlook the absence of fine-grained waveform features in ECG reports. Additionally, due to medical LLM hallucinations Huang et al. (2023); Günay et al. (2024) and the variability of waveform features for the same disease across patients Jin (2018), relying solely on LLMs Yu et al. (2024) for report augmentation is unreliable. To address these challenges, we propose the FG-CLEP training process.
False Negatives in Contrastive Learning Traditional multi-modal contrastive learning Radford et al. (2021) assumes that only images and captions from the same record are positive pairs. However, this assumption often fails in the ECG domain, where most ECGs are normal, and abnormalities typically involve common diseases, leading to frequent false negatives. There have been several attempts to address this issue Lavoie et al. (2024); Jiang et al. (2023b); Sun et al. (2023); Li et al. (2023); Wang et al. (2022). Some approaches Jiang et al. (2023b); Li et al. (2023) attempt to add a regularization term to mitigate false negatives. Others Sun et al. (2023); Wang et al. (2022) introduce a matrix to measure the similarity between different reports, guiding contrastive learning to identify and address false negatives. In this paper, we explore the application of the latter approach in the ECG multi-modal contrastive learning domain.
3 Method
In this section, we present the technical details of the FG-CLEP framework, which includes a novel pre-training pipeline that recovers missing waveform features in reports and helps the model learn fine-grained ECG representations, as illustrated in Figure 2, as well as a specific model architecture tailored for ECG, shown in Figure 3. This architecture is designed to address the frequent false negatives in ECG-Report pairs and the multi-label nature of ECG downstream tasks. It consists of the following components: (1) ECG and text encoders that extract embeddings, (2) a Semantic Similarity Matrix that measures the similarity of reports from different patients, and (3) $L_{sig}$ and $L_{fnm}$ that train the entire model.
3.1 Model Architecture
As illustrated in Figure 3, FG-CLEP consists of one ECG encoder and one text encoder.
ECG Encoder. We encode ECG signals into embeddings $\mathbf{e}\in\mathbb{R}^{D}$ using an ECG encoder $E_{\text{ecg}}$. A projection head then maps raw embeddings to $\mathbf{e}_{p}\in\mathbb{R}^{P}$.
$\mathbf{e}$ $=E_{\text{ecg}}(x_{\text{ecg}})$ (1a)
$\mathbf{e}_{p}$ $=f_{e}(\mathbf{e})$ (1b)
where $f_{e}$ is the projection head of the ECG encoder.
Text Encoder We create clinically meaningful text embeddings $\mathbf{t}\in\mathbb{R}^{M}$ by a text encoder. We project them to $\mathbf{t}_{p}\in\mathbb{R}^{P}$ as
$\mathbf{t}$ $=E_{\text{txt}}(x_{\text{txt}})$ (2a)
$\mathbf{t}_{p}$ $=f_{t}(\mathbf{t})$ (2b)
where $f_{t}$ is the projection head and $E_{\text{txt}}$ denotes the text encoder. This gives the same embedding dimension $P$ as the ECG encoder, suitable for contrastive learning.
Semantic Similarity Matrix False negatives in the pre-training phase arise from the assumption that ECGs and reports from different patients are unmatched. However, due to the prevalence of common diagnoses, many ECGs exhibit similar symptoms, leading to reports with similar semantics. To address this, we introduce a Semantic Similarity Matrix similar to Sun et al. (2023); Wang et al. (2022) to measure the similarity of reports from different patients during the pre-training phase.
We denote an ECG-report dataset as $D=\{(x_{\text{ecg}_{i}},x_{\text{txt}_{i}})\mid i\in[0,n)\}$, where $(x_{\text{ecg}_{i}},x_{\text{txt}_{i}})$ represents a sample with paired ECG-report content. The ECG and text signals are encoded into $(\mathbf{e}_{p},\mathbf{t}_{p})$ as disscussed above, and the semantic similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ is defined as follows:
$\mathbf{S}_{ij}=\text{sim}(\mathbf{t}_{pi},\mathbf{t}_{pj})$
where $\text{sim}(\cdot,\cdot)$ is the cosine similarity function.
$\text{sim}(\mathbf{t}_{pi},\mathbf{t}_{pj})=\frac{\mathbf{t}_{pi}\cdot\mathbf{t}_{pj}}{\|\mathbf{t}_{pi}\|\|\mathbf{t}_{pj}\|}$
This results in a similarity matrix $\mathbf{S}$, where each element $\mathbf{S}_{ij}$ indicates the similarity between the $i$-th and $j$-th report embeddings. The Semantic Similarity Matrix guides the contrastive learning process and reduces false negatives by considering the semantic similarities across different pairs. We give an intuitive visualization of the semantic similarity matrix in Section 5.1.
Loss Function The loss function of our FG-CLEP framework consists of two components: $L_{\text{sig}}$ and $L_{\text{fnm}}$. $L_{\text{sig}}$ represents a sigmoid-based loss function tailored to adapt to downstream ECG multi-label classification tasks. $L_{\text{fnm}}$ addresses the issue of false negatives commonly observed in ECG datasets due to the prevalence of common diagnoses.
Sigmoid-based Contrastive Loss $L_{\text{sig}}$ Traditional multimodal contrastive learning often relies on the InfoNCE loss Oord et al. (2018), which involves a softmax operation. However, for ECG downstream tasks, which are mostly multi-label classification problems, using the softmax operation is suboptimcal. To address this problem, we propose using a sigmoid-based loss during the pre-training phase. Specifically, we leverage the sigmoid loss Zhai et al. (2023), which is known for its memory efficiency compared to InfoNCE loss Oord et al. (2018). Here, we use it for its disuse of the softmax operation, making it more suitable for downstream ECG multi-label classification tasks.
The $L_{\text{sig}}$ is defined as:
$L_{\text{sig}}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\log\left(\frac{1}{1+e^{-z_{ij}(-t\cdot\text{sim}(\mathbf{e}_{pj},\mathbf{t}_{pj})+b)}}\right)$
Where $B$ represents the batch size, $z_{ij}$ denotes the match between a given ECG and report input, equaling 1 if they are paired and -1 otherwise. At initialization, the significant imbalance caused by numerous negatives dominates the loss, resulting in large initial optimization steps aimed at correcting this bias. To mitigate this issue, an additional learnable bias term $b$, similar to the temperature $t$, is used. This ensures that the training starts approximately close to the prior, avoiding the need for substantial over-correction.
Figure 3: The Overview of Model Architecture.
False Negative Mitigation Loss $L_{\text{fnm}}$ Due to the long-tail distribution of ECG labels, most electrocardiograms have similar reports. Simply aligning the embeddings of paired electrocardiograms with their corresponding reports while distinguishing them from unpaired reports results in false negatives. To address this issue, we incorporate the semantic similarity matrix $S$ obtained above into the loss function to guide contrastive learning, similar to Sun et al. (2023). This matrix captures the semantic similarity of reports from different patients, allowing us to identify and correct false negative samples.
The $L_{\text{fnm}}$ is defined as:
$L_{\text{fnm}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\left|\text{sim}(\mathbf{e}_{pi},\mathbf{t}_{pj})-\mathbf{S}_{ij}\right|$
where $B$ represents the batch size, $\mathbf{S}_{ij}$ is the semantic similarity between the $i$-th and $j$-th report embeddings, and the ||$\cdot$|| ⋅ | represents the L1-distance. We alleviates the false negative problem by reducing the L1 distance between $\text{sim}(\mathbf{e}_{pi},\mathbf{t}_{pj})$ and $\mathbf{S}_{ij}$.
Combined Loss Function The final loss function combines $L_{\text{sig}}$ and $L_{\text{fnm}}$:
$L=L_{\text{sig}}+\lambda L_{\text{fnm}}$
where $\lambda$ is a hyperparameter that controls the trade-off between the two loss components.
In conclusion, by integrating the semantic similarity matrix $\mathbf{S}$ and using a sigmoid-based contrastive loss, our pretraining loss function mitigates the frequent false negatives and better adapts to downstream ECG multi-label tasks when performing zero-shot prediction, leading to improved learning of meaningful representations of ECG signals.
Inference During the inference phase, given an ECG and a text description, we calculate their similarity as described above and use the sigmoid function to determine the probability. The probability score $P$ for the matching between the ECG embedding $\mathbf{e}_{p}$ and the text embedding $\mathbf{t}_{p}$ is computed as follows:
$P=\sigma(\text{sim}(\mathbf{e}_{p},\mathbf{t}_{p}))$
where $\sigma$ denotes the sigmoid function. This probability score indicates how likely it is that the ECG and the text description are related. The higher the score, the more confident we are in the match.
3.2 Training Process
As discussed above, waveform features, which serve as intermediate products of diagnostic reasoning by doctors, are often not recorded in reports but are crucial for the model to learn ECG representations and understand the diagnostic process. Therefore, we propose a novel training process illustrated in Figure 2 consisting of three steps: (1) training the CLEP model using contrastive learning on original ECG-report pairs, (2) generating potential waveform features based on the original report using LLMs and validating them with CLEP, and (3) continuing to train the CLEP model with this augmented report containing waveform features to obtain the final FG-CLEP model.
The key of our training process is to recover the missing waveform features in the report. Given a report, we query LLMs with the question, ‘What waveform features are most likely to be present in electrocardiograms with these symptoms?’ to identify potential overlooked waveform features. To format the results, we further instruct, ‘Organize these waveform features into a Python list, with each item representing a distinct waveform feature.’ Using this explicit chain-of-thought instruction Wei et al. (2022), we obtain a list of potential waveform features. However, LLM outputs are unreliable for two reasons: First, ECG waveform features and diagnoses are not in a one-to-one correspondence Jin (2018)—a single disease may present different waveform characteristics across individuals. Doctors can infer a diagnosis from waveform features, but not the other way around. Second, even if the non-bijective relationship is excluded, the LLM’s output is inherently unstable due to hallucination issues Huang et al. (2023); Günay et al. (2024).
Thus, we validate LLM-generated waveform features by computing their similarity to the ECG signal using coarse CLEP, selecting only high-confidence waveform features for augmentation. These validated features are then incorporated into the original report for continued training of the Fine-Grained CLEP model. See Appendix B for pseudo code.
Table 1: Results of zero-shot classification. ENS: ensemble inference.
macro AUC PTB-XL-Super PTBXL-Sub PTBXL-Form PTBXL-Rhythm CPSC2018 CSN
METS Li et al. (2024) 76.31 80.12 65.95 86.29 82.49 77.20
FG-METS 78.12$\uparrow$1.81 82.01$\uparrow$1.89 66.33$\uparrow$0.38 90.12$\uparrow$3.83 86.92$\uparrow$4.43 81.20$\uparrow$4.00
MERL Liu et al. (2024b) 74.20 75.70 65.90 78.50 82.80 74.40
FG-MERL 76.70$\uparrow$2.50 78.20$\uparrow$2.50 66.80$\uparrow$0.90 81.00$\uparrow$2.50 85.30$\uparrow$2.50 76.90$\uparrow$2.50
CLEP 77.50 81.85 66.29 88.60 85.15 80.10
FG-CLEP 79.28$\uparrow$1.78 83.57$\uparrow$1.72 67.77$\uparrow$1.48 92.31$\uparrow$3.71 88.24$\uparrow$3.09 82.46$\uparrow$2.36
CLEPENS 75.64$\downarrow$1.86 82.55$\uparrow$0.70 64.74$\downarrow$1.55 88.67$\uparrow$0.07 83.91$\downarrow$1.24 81.24$\uparrow$1.14
FG-CLEP ENS 79.68$\uparrow$0.40 83.65$\uparrow$0.08 70.79$\uparrow$3.02 91.52$\downarrow$0.79 87.08$\downarrow$1.16 84.60$\uparrow$2.14
4 Experiments
4.1 Datasets
We pre-train the FG-CLEP framework using the MIMIC-ECG Gow et al. dataset and test it on the PTB-XL Wagner et al. (2020), CPSC2018 Liu et al. (2018), and CSN Zheng et al. (2022) datasets, following the benchmark proposed by Liu et al. (2024b). All the ECGs in the datasets are 12-lead recordings. The MIMIC-ECG dataset contains nearly 800,000 ECG-report pairs. To improve data quality, we excluded samples with an empty report or reports containing fewer than three words, removed reports without useful information, and discarded ECGs with unexpected situations. Details regarding the train:validation:test split and other dataset-specific information are provided in Appendix A.
Table 2: Results of Linear Evaluation.
Method PTB-XL-Super PTBXL-Sub PTBXL-Form PTBXL-Rhythm CPSC2018 CSN
1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100%
Random Init 70.45 77.09 81.61 55.82 67.60 77.91 55.82 62.54 73.00 46.26 62.36 79.29 54.96 71.47 78.33 47.22 63.17 73.13
SimCLRChen et al. (2020) 63.41 69.77 73.53 60.84 68.27 73.39 54.98 56.97 62.52 51.41 69.44 77.73 59.78 68.52 76.54 59.02 67.26 73.20
BYOLGrill et al. (2020) 71.70 73.83 76.45 57.16 67.44 71.64 48.73 61.63 70.82 41.99 74.40 77.17 60.88 74.42 78.75 54.20 71.92 74.69
BarlowTwinsZbontar et al. (2021) 72.87 75.96 78.41 62.57 70.84 74.34 52.12 60.39 66.14 50.12 73.54 77.62 55.12 72.75 78.39 60.72 71.64 77.43
MoCo-v3Chen et al. (2021) 73.19 76.65 78.26 55.88 69.21 76.69 50.32 63.71 71.31 51.38 71.66 74.33 62.13 76.74 75.29 54.61 74.26 77.68
SimSiamChen and He (2021) 73.15 72.70 75.63 62.52 69.31 76.38 55.16 62.91 71.31 49.30 69.47 75.92 58.35 72.89 75.31 58.25 68.61 77.41
TS-TCCEldele et al. (2021) 70.73 75.88 78.91 53.54 66.98 77.87 48.04 61.79 71.18 43.34 69.48 78.23 57.07 73.62 78.72 55.26 68.48 76.79
CLOCSKiyasseh et al. (2021) 68.94 73.36 76.31 57.94 72.55 76.24 51.97 57.96 72.65 47.19 71.88 76.31 59.59 77.78 77.49 54.38 71.93 76.13
ASTCLWang et al. (2023) 72.51 77.31 81.02 61.86 68.77 76.51 44.14 60.93 66.99 52.38 71.98 76.05 57.90 77.01 79.51 56.40 70.87 75.79
CRTZhang et al. (2023) 69.68 78.24 77.24 61.98 70.82 78.67 46.41 59.49 68.73 47.44 73.52 74.41 58.01 76.43 82.03 56.21 73.70 78.80
ST-MEMNa et al. (2024) 61.12 66.87 71.36 54.12 57.86 63.59 55.71 59.99 66.07 51.12 65.44 74.85 56.69 63.32 70.39 59.77 66.87 71.36
MERLLiu et al. (2024b) 82.39 86.27 88.67 64.90 80.56 84.72 58.26 72.43 79.65 53.33 82.88 88.34 70.33 85.32 90.57 66.60 82.74 87.95
CLEP 84.04 88.79 89.82 69.09 86.08 92.50 67.89 72.35 82.59 61.79 91.86 90.18 83.12 93.42 96.56 63.00 80.03 93.35
FG-CLEP 84.89 89.51 90.77 69.96 85.75 92.62 68.91 74.80 85.42 68.99 91.35 94.08 83.35 93.60 96.65 62.59 79.35 93.46
4.2 Implementation Details
Pre-training Implementation: In the pre-training stage, we utilize a randomly initialized 1D-ResNet50 model He et al. (2016) as the ECG encoder and BioClinicalBERT Alsentzer et al. (2019) for text encoding. The AdamW optimizer is selected with a learning rate of $2\times 10^{-5}$ and a weight decay of $1\times 10^{-4}$. CLEP is pre-trained for 10 epochs with original reports and FG-CLEP is trained for another 3 epochs with fine-grained reports, using a cosine annealing scheduler for learning rate adjustments and a warmup phase for the first 10% of training steps. A batch size of 100 is maintained. The temperature parameters $t$ and $b$ are initialized to $\log 10$ and $-10$, respectively. The default hyperparameter $\lambda$ is set to 0.5 and the default threshold for selecting high-confidence waveform features is set to 0.95. We use LLaMA3-8B AI@Meta (2024) as our LLM to query potential waveform features and use vLLM Kwon et al. (2023) to speed up inference. All experiments used two NVIDIA A800 80GB GPUs, except LLaMA3-70B ablation, which used four.
Downstream Task Implementation: We evaluated the downstream tasks using both zero-shot and linear probe settings. For the zero-shot setting, we froze the entire model and used the text of each category as the prompt. We computed the similarity between the ECG embedding and the category text embedding as the classification probability. Additionally, we employed an ensemble method to enhance zero-shot performance. Specifically, in addition to using the category as text, we also added ‘category in lead x’ (x represents any of the 12 leads) as text to compute the probability and used the highest probability as the final probability for that category. For linear probing, we kept the ECG encoder frozen and updated only the parameters of a newly initialized linear classifier. We conducted linear probing for each task using 1%, 10%, and 100% of the training data. For all downstream tasks, we used macro AUC as the metric.
4.3 Zero-Shot Ability
The zero-shot results are illustrated in Table 1. Both CLEP and FG-CLEP performed well. A detailed examination of the data reveals that FG-CLEP significantly outperforms CLEP on PTBXL-Form, PTBXL-Rhythm demonstrating that continue training using fine-grained reports substantially enhanced the model’s ability to capture local ECG waveform features. This improvement is particularly evident when using the ensemble method, which extends the label text to 12 leads (‘label in lead x’, where x represents any of the 12 leads). This further indicates FG-CLEP’s fine-grained waveform feature capture capability. However, the ensemble inference method often proves detrimental to CLEP, as seen in PTBXL-Super, PTBXL-Form, and CPSC2018. Additionally, we applied our generated fine-grained reports to other methods, METS Li et al. (2024) and MERL Liu et al. (2024b), to validate the generalizability. The results demonstrate that our fine-grained reports can also enhance the performance of these methods.
4.4 Linear Evaluation
We aim to evaluate the learned model transferability to downstream supervised tasks. We froze the ECG encoder and fine-tuned a randomly initialized linear classification head on the training data with binary cross-entropy loss. We compared a series of contrastive and generative self-supervised learning methods. Results in Table 2 show that FG-CLEP still achieves the best performances across all methods in most scenarios.
Furthermore, when comparing the linear probe result in Table 2 with the zero-shot result in Table 1, we surprisingly find that FG-CLEP’s zero-shot predictions are comparable to Linear Probe evaluations using 10% of the data in PTBXL-Sub, PTBXL-Form, CPSC2018, and CSN. Additionally, the zero-shot performance in PTBXL-Form is comparable to the full 100% Linear Probe evaluation. This further confirms the robustness and generalizability of our framework.
5 Analysis
In this section, we visualize the semantic similarity matrix to identify false negatives, conduct an ablation study to validate FG-CLEP’s effectiveness, and assess its performance across various LLMs, text encoders, and ECG encoders. The reported metrics reflect the average zero-shot AUC across the aforementioned six datasets.
5.1 Semantic Similarity Matrix
We visualize the semantic similarity matrix in Figure 4. The left side shows the semantic similarity matrix from a random batch. As illustrated, ECGs and reports from different records may share similarities to some extent. Ignoring these similarities would result in a diagonal matrix with ones on the diagonal and zeros elsewhere, which is obviously wrong. The right side displays a semantic similarity matrix where the first 16 entries are normal ECGs and the last 16 are abnormal ECGs. The matrix effectively captures the semantic similarities of the normal ECGs.
Figure 4: The Heatmap of Semantic Similarity Matrix. Left: from a random batch; Right: with the first 16 as normal ECG and the last 16 as abnormal ECG.
5.2 Ablation Study
To evaluate the effectiveness of our proposed fine-grained pretraining strategy, we conducted an ablation study, with the results presented in Table 3. Specifically, we compared FG-CLEP with CLEP trained without the additional 3 epochs of fine-grained report training. The results demonstrate that FG-CLEP significantly outperforms CLEP. To further confirm that the performance gains stem from the fine-grained reports rather than additional training epochs, we trained CLEP for 3 extra epochs using the original reports. The findings indicate that CLEP converges within 10 epochs, and the additional training even risks overfitting, leading to a slight performance drop. This further validates the importance of fine-grained reports.
We also tested training FG-CLEP from scratch using fine-grained reports instead of continuing from CLEP. The results show no performance improvement but increased computational costs, supporting our default approach of continuing training from CLEP. Additionally, to verify the effectiveness of our model architecture in mitigating false negatives, we evaluated the performance using the original InfoNCE loss. The results reveal a significant performance drop, highlighting the efficacy of our proposed loss function in addressing false negatives.
Table 3: Results of ablation study.
Model Setting AUC
FG-CLEP (continued from CLEP) 82.27
FG-CLEP (trained from scratch) 82.28
CLEP (without fine-grained training) 79.92
CLEP (+3 epochs, original reports) 79.01
CLEP (InfoNCE loss only) 80.78
Finally, to assess the robustness of our method, we conducted ablation experiments on the loss hyperparameter $\lambda$ and the threshold for selecting fine-grained waveform features. As shown in Figure 5, our model maintains strong performance across different $\lambda$ and threshold values, demonstrating its robustness. To ensure high precision in generating waveform features, we set a relatively high default threshold of 0.95.
Figure 5: Effect of Threshold and $\lambda$
5.3 Different Components
We conducted experiments to evaluate the performance of our framework using different LLMs, text encoders, and ECG encoders. The results are presented in Table 4. The findings indicate that our framework is robust across various components. Specifically, for different LLMs Abdin et al. (2024); Jiang et al. (2023a); AI@Meta (2024); Labrak et al. (2024); Ankit Pal (2024), larger LLM provide some performance improvements, though the gains are not substantial. While domain-specific LLMs possess more medical knowledge, our method requires formatting the waveform feature outputs, an area where domain-specific models are less effective, resulting in performance that does not surpass general-purpose models. For different text encoders Alsentzer et al. (2019); Gu et al. (2021); Jin et al. (2023); Lee et al. (2020), our framework consistently achieves significant improvements. Regarding ECG encoders He et al. (2016); Dosovitskiy et al. (2020), scaling up the model size leads to noticeable performance gains, and ResNet He et al. (2016) outperforms ViT Dosovitskiy et al. (2020) in capturing ECG features.
Table 4: Results on Different LLM/Text Encoder/ECG Encoder.
Model Type CLEP FG-CLEP $\uparrow$
Different LLMs
Phi-3-mini-4k-instruct 79.92 81.51
Mistral-7B-Instruct-v0.2 79.92 81.89
LLaMA3-8B-Instruct 79.92 82.27
LLaMA3-70B-Instruct 79.92 82.80
BioMistral-7B 79.92 81.67
LLaMA3-OpenBioLLM-8B 79.92 82.36
Different Text Encoders
BioClinicalBERT 79.92 82.27
PubMedBERT 80.10 81.87
Med-CPT 78.25 80.87
BioBERT 78.33 80.20
Different ECG Encoders
ResNet18 79.70 81.26
ResNet50 79.92 82.27
ResNet152 80.35 82.98
ViT-Base 76.20 78.12
ViT-Large 76.14 78.04
ViT-Huge 76.31 78.28
6 Conclusion
In this paper, we introduced FG-CLEP, a novel approach for fine-grained contrastive learning in ECG-text tasks, addressing the critical issue of incomplete ECG reports, particularly the absence of key waveform features. By incorporating LLMs to generate potential waveform features and validating them with our pipeline, we were able to enhance the quality of ECG reports and better capture the diagnostic reasoning process. Additionally, we tackled the challenges associated with multi-label classification and frequent false negatives by introducing a sigmoid-based loss function and a semantic similarity matrix to guide contrastive learning respectively. Experimental results across six ECG datasets, including PTB-XL, CPSC2018, and CSN, demonstrate that FG-CLEP outperforms existing state-of-the-art methods in both zero-shot prediction and linear probing, highlighting its effectiveness in improving ECG classification and facilitating more accurate diagnostic insights.
Limitations
Although our method effectively leverages large language models (LLMs) to recover crucial waveform features from ECG reports under the challenges of hallucinations and the non-bijective relationship between waveform features and diagnoses, there are certain limitations. One of the main drawbacks is the computational cost and time required, especially during the pre-training phase. The volume of ECG data is extensive, and generating accurate waveform features for such large datasets using LLMs is both time-consuming and resource-intensive. Future research should focus on developing more efficient methods for recovering waveform features.
Appendix A Dataset Analysis
We pre-train the FG-CLEP using the MIMIC-ECG dataset and test it on the PTB-XL, CPSC2018, and CSN datasets. All the ECGs in the datasets are 12-lead recordings. The PTB-XL dataset can be further divided into four subsets, and we follow the official train:validation:test split. For CPSC2018 and CSN, we split the dataset as 70%:10%:20% for the train:validation:test split. The statistics of the datasets used are presented in Table 5.
MIMIC-ECG The MIMIC-ECG dataset contains nearly 800,000 ECG-report pairs from approximately 160,000 unique patients. These diagnostic ECGs utilize 12 leads and are 10 seconds in duration, with a sampling rate of 500 Hz. https://physionet.org/content/mimic-iv-ecg/1.0/.
PTB-XL The PTB-XL ECG dataset is a large dataset of 21,799 clinical 12-lead ECGs from 18,869 patients of 10-second length. There are four subsets with multi-label classification tasks: Superclass (5 categories), Subclass (23 categories), Form (19 categories), and Rhythm (12 categories). Notably, these four subsets have different numbers of samples. https://physionet.org/content/ptb-xl/1.0.3/.
CPSC2018 This publicly accessible dataset comprises 6,877 standard 12-lead ECG records, each sampled at a rate of 500 Hz, with durations ranging from 6 to 60 seconds. The dataset is annotated with 9 distinct labels. http://2018.icbeb.org/Challenge.html.
Chapman-Shaoxing-Ningbo (CSN) This dataset contains 12-lead ECGs of 45,152 patients with a 500 Hz sampling rate. It features multiple common rhythms and additional cardiovascular conditions, all labeled by professional experts. https://physionet.org/content/ecg-arrhythmia/1.0.0/.
Table 5: The statistics of used datasets.
Pretrain # ECGs # Reports
MIMIC-ECG 773,268 773,268
Evaluation # Train # Valid # Test # Classes
PTB-XL Super 17,084 2,146 2,158 5
PTB-XL Sub 17,084 2,146 2,158 23
PTB-XL Form 7,197 901 880 19
PTB-XL Rhythm 16,832 2,100 2,098 12
CPSC2018 4,800 684 1,383 9
CSN 31,606 4,515 9,031 51
Appendix B Pseudo Code
The pseudo-code of our FG-CLEP training process is shown in algorithm 1
Algorithm 1 FG-CLEP Training Process
1:  Input: $D=\{(x_{\text{ecg}_{i}},x_{\text{txt}_{i}})\mid i\in[0,n)\}$
2:  Output: FG-CLEP
3:  Perform contrastive training on CLEP using $D$
4:  Generate fine-grained reports
5:  for $i=0\text{ to }n-1$ do
6:     $f_{\text{features}}=\text{LLM}(x_{\text{txt}_{i}},\text{prompt})$
7:     for $j=1\text{ to }m$ do
8:        where $m$ is the number of waveform features generated
9:        $\text{sim}=\text{CLEP}(x_{\text{ecg}_{i}},f_{j})$
[…]
11:           $x_{\text{txt}_{i}}=x_{\text{txt}_{i}}+f_{j}$
12:        end if
13:     end for
14:  end for
15:  Continue training CLEP on $\{(x_{\text{ecg}_{i}},x_{\text{txt}_{i}})\}$ to obtain FG-CLEP
Appendix C Embedding Visualization
We also demonstrate the effectiveness of our representation learning framework by plotting t-SNE visualizations of ECG embeddings produced for PTB-XL ECGs in five classical waveform features. As shown in Fig. 6, our model produces well-clustered representations. Furthermore, as expected, FG-CLEP learns more fine-grained local waveform features of ECGs. Specifically, FG-CLEP clusters ‘prolonged PR interval’ much better than CLEP.
Figure 6: Embeddings visualization of PTB-XL ECGs in 5 waveform features by (a)CLEP and (b)FG-CLEP. Dimension reduced by t-SNE.
Appendix D Running Cases for Generating Fine-Grained Reports
We present three case studies illustrating how the fine-grained reports with waveform features are generated step by step, as shown in Figure 7, 8, 9. The LLM is prompted with two explicit chain-of-thought instructions: ‘What waveform features are most likely to be present in electrocardiograms with these symptoms?’ followed by ‘Organize these waveform features into a Python list, with each item representing a distinct waveform feature.’
Figure 7: Running Case1 with Atrial fibrillation and Right bundle branch block.
Figure 8: Running Case2 with Myocardial Infarction.
Figure 9: Running Case3 with Hypertrophy.
Appendix E ECG-Text Retrieval
We attempted to use FG-CLEP to retrieve electrocardiograms (ECGs) from the MIMIC-ECG dataset through text. To test our model’s ability to capture fine-grained waveform features, we tested a series of typical waveform features such as ‘RSR’ Pattern,’ ‘Inverted T-waves,’ and ‘Low QRS voltages.’ Figure 10 shows the Top 3 retrieved ECGs with probabilities all greater than 0.99. Our model demonstrated strong capability in retrieving ECGs through waveform feature text, which can lead to two applications: (1) Helping doctors quickly find similar cases or specific ECG patterns, aiding in diagnosis and treatment decision-making; (2) In medical education, text-based retrieval can quickly find typical ECG cases, assisting in teaching and training, thereby improving educational effectiveness.
Figure 10: Top 3 retrieved ECG using FG-CLEP.